OpenStudy (anonymous):

A funnel with a height of 3.42 inches is shown below. FLVS_GeomB_v13_08.06_G7Q1 The funnel has a diameter of 1.56 inches that includes a wall that is 0.125 inches thick. What is the maximum volume of the funnel rounded to the nearest hundredth?

OpenStudy (callisto):

if the width of wall is not a concern, max volume = 1/3 pi [(1.56-0.125)/2]^2 x 3.42 = 1.844 in^3 if the width of wall is a concern, max volume = 1/3 pi [(1.56-0.125)/2]^2 x (3.42-0.125) = 1.776 in^3